# Large Language Model Confidence Estimation via Black-Box Access

### v3

@@ 1 Introduction
We observe in the experiments that we outperform state-of-art baselines for black-box LLM confidence estimation on standard metrics such as Area Under the Receiver Operator Characteristic (AUROC) and Area Under Accuracy-Rejection Curve (AUARC), where improvements in AUROC are over $10\%$ in some cases. The confidence model being interpretable we also analyze which features are important for different LLM and dataset combinations. We interestingly find that for a given dataset the important features are shared across LLMs. Intrigued by this finding we apply confidence models built for one LLM to the responses of another and further find that they generalize well across LLMs. This opens up the possibility of simply building a single (universal) confidence model for some chosen LLM and zero shot applying it to other LLMs on a dataset.
@@ 3 Methodology
We now describe our methodology to estimate confidences for individual LLM outputs.
@@ 3.1 Elicitation of Variable LLM Behavior
We first propose six black-box strategies that can elicit variable behavior in an LLM indicative of how trustworthy its output is likely to be. Based on this variability we construct features for our confidence model in the next subsection. Note that all strategies may not be relevant in all cases. For instance, some of the strategies require a context in the prompt, while others such as SRC require longer responses (two or more sentences). For all the perturbations but for Stochastic Decoding and Split Response Consistency, the perturbations are applied to the context if available or to the question of the input.
@@
Stochastic Decoding (SD): This is the simplest strategy which is also done in previous works. Here the prompt is not varied, but rather using various decoding strategies comprising of greedy , beam search and nucleus sampling [11] multiple outputs are sampled. As seen in Table 1 first row after sampling one could have four different outputs, which could be indicative of the LLM not being confident in its response. Specifically in the experiments, we obtain one generation using greedy and beam search decoding technique and 3 generations using nucleus sampling.
@@
Paraphrasing (PP): In this strategy we paraphrase the context in the prompt and observe how that changes the output. An example of this is shown in Table 1. For paraphrasing, we use back translation, where we convert the original prompt into another language and translate it back into English. We use machine translation models from Helsinki-NLP on huggingface and translate the text from English to French and then back to English. This new prompt then can be used to query the LLM. Changes to the output could indicate brittleness in the LLMs original response. One could also prompt an LLM to paraphrase the responses, however, in our initial experiments, we observed that when context is involved, the model does not paraphrase the entire context and parts of it were omitted.
@@
Sentence Permutation (SP): If the input has several named entities, we noticed that when the order of the named entities is changed without changing the meaning of the sentence, the output of the LLM also varied. We first use named entity detector to identify the named entities and then randomly reorder certain number of these sentences. An example of this is seen in Table 1 third row, where the last sentence in the prompt is now the second sentence. As such, if the number of sentences with named entities is less than five, we reorder all of them. If it is greater than five, then we randomly select five and reorder them. Most such reorderings do not affect the LLM output if it is confident.
@@
Entity Frequency Amplification (EFA): Similar to above, repeating sentences with named entities could also throw off the model’s outputs. We sample a sentence from all the sentences with named entities and repeat it three times. Again, here too the output of the LLM should be maintained if the LLM is confident. An example of this is seen in Table 1 fourth row, where the first sentence is repeated twice.
@@
Stopword Removal (SR): We remove stopwords from the context as specified by the NLTK library. Stopwords are commonly occurring words (viz. "the", "are", "to", etc.) that are assumed to have limited context specific information. Removal of such words should ideally not alter the response of an LLM if the LLM is certain of the answer. An example of this is seen in Table 1 fifth row, where the stopwords are striked out. We ensured that the negative words were not removed as they would change the meaning of the sentence.
@@
Split Response Consistency (SRC): In this case like in the SD case the prompt is not perturbed. Rather the output is analyzed where it is randomly split such that each part is at least a single sentence. Semantic inconsistency between the two parts is measured using an NLI models contradiction probability, where one part is taken as the premise and the other the hypothesis. An example of this is seen in Table 1 last row, where the two sentences are clearly at odds with each other. This strategy though requires that the response is at least a couple of sentences long.
@@
As seen in Table 2, the four perturbations above (PP, SP, EFA and SR) that alter the original prompt still maintain the semantics as intended in almost all cases.
$99.81$% $99.23$% $99.66$% $99.12$%
@@ 3.2 Featurization
SRC Minimum Value: As mentioned above, semantic inconsistency between the two parts is measured using an NLI models contradiction probability, where one part is taken as the premise and the other the hypothesis. The highest contradiction probability amongst multiple such partitions is the feature value for this strategy. In Table 1 last row, there are only two sentences so only one split would be done and since the sentences contradict each other the NLI contradiction probability would be high or consistency would be low.
Note that optionally one can also pass the entire prompt as a feature in addition to the above. In the experiments, we saw minimal improvement with such an addition. Semantic set and lexical similarity were first used by [18] where they applied it only for SD perturbation discussed in the previous section.

### v4

@@ 1 Introduction
We observe in the experiments that we outperform state-of-art baselines for black-box LLM confidence estimation on standard metrics such as Area Under the Receiver Operator Characteristic (AUROC), Area Under Accuracy-Rejection Curve (AUARC) and Expected Calibration Error (ECE), where improvements in AUROC are over $10\%$ in some cases. The confidence model being interpretable we also analyze which features are important for different LLM and dataset combinations. We interestingly find that for a given dataset the important features are shared across LLMs. Intrigued by this finding we apply confidence models built for one LLM to the responses of another and further find that they generalize well across LLMs. This opens up the possibility of simply building a single (universal) confidence model for some chosen LLM and zero shot applying it to other LLMs on a dataset.
@@ 3.1 Elicitation of Variable LLM Behavior
The baselines generated multiple responses from the model and extracted features from these generations to quantify how much the responses diverge. This divergence metric is then used to estimate the confidence of the model. We conjecture that rather than just eliciting multiple generations using the same input, perturbing the input and then generating a myriad of responses would provide a better estimate of the model’s confidence. To that end, we use six perturbations to elicit the model’s response. We first propose six black-box strategies that can elicit variable behavior in an LLM indicative of how trustworthy its output is likely to be. Based on this variability we construct features for our confidence model in the next subsection. Note that all strategies may not be relevant in all cases. For instance, some of the strategies require a context in the prompt, while others such as SRC require longer responses (two or more sentences). For all the perturbations but for Stochastic Decoding and Split Response Consistency, the perturbations are applied to the context if available or to the question of the input.
@@
Stochastic Decoding (SD): Similar to previous works, we sample the LLM to obtain various generations. The Algorithm 1 elaborates it further. As seen in Table 1 first row after sampling one could have four different unique outputs, which could be indicative of the LLM not being confident in its response.
$\begin{array}[]{rl}output&\;\leftarrow\;\texttt{generate one response from themodel using greedy decoding}\\ &+\texttt{one response using beam search decoding}\\ &+\texttt{three responses using nucleus sampling}\end{array}$
Algorithm 1 Stochastic Decoding (SD): Algorithm to collect decoded features under various strategies.
@@
Paraphrasing (PP): In this strategy, we paraphrase the context in the prompt and observe how that changes the output. An example of this is shown in Table 1. It is illustrated in listing 2 For paraphrasing, we use back translation, where we convert the original prompt into another language and translate it back into English. We use machine translation models from Helsinki-NLP (referred to as MT-Model in the Algorithm) on Huggingface. One could also prompt an LLM to paraphrase the responses, however, in our initial experiments, we observed that when context is involved, the model does not paraphrase the entire context and parts of it were omitted.
$\begin{array}[]{rl}translated\_sentence&\;\leftarrow\;\texttt{Translate from English to French}\\ back\_translated\_sentence&\;\leftarrow\;\texttt{Translate the above sentence from French to English}\end{array}$
Algorithm 2 Paraphrasing (PP): Algorithm for paraphrasing via back-translation.
@@
Sentence Permutation (SP): If the input has several named entities, we noticed that when the order of the named entities is changed without changing the meaning of the sentence, the output of the LLM also varied. We explain it in algorithm 3 We first use a Named Entity Recognizer (NER) to identify the named entities and then randomly reorder these sentences. An example of this is seen in Table 1 third row, where the last sentence in the prompt is now the second sentence. We use NLTK for identifying atomic sentences and spacy for NER.
$\begin{array}[]{rl}sentences&\;\leftarrow\;\texttt{split sentences}\\ tagged\_sentences&\;\leftarrow\;\texttt{tag named entities using NER model}\\ NER\_sentences&\;\leftarrow\;\texttt{Sample 5 sentences with named entities}\\ sentences&\;\leftarrow\;\texttt{remove the NER\_sentences from sentences }\\ \texttt{for sentence in NER\_sentences:}\\ final\_sentences&\;\leftarrow\;\texttt{Insert sentence randomly into the {sentences} }\end{array}$
Algorithm 3 Sentence Permutation (SP): Algorithm for random sentence permutation via NER sampling.
@@
Entity Frequency Amplification (EFA): Similar to above, repeating sentences with named entities could also throw off the model’s outputs. Again, here too the output of the LLM should be maintained if the LLM is confident. An example of this is seen in Table 1 fourth row, where the first sentence is repeated twice. We use NLTK for identifying atomic sentences and spacy for NER.
$\begin{array}[]{rl}sentences&\;\leftarrow\;\texttt{split sentences}\\ tagged\_sentences&\;\leftarrow\;\texttt{NER\_model tags the sentences}\\ NER\_sentence&\;\leftarrow\;\texttt{Sample one sentence from the tagged\_sentences}\\ sentences&\;\leftarrow\;\texttt{repeat the NER\_sentence two more times in thesentences array}\\ \end{array}$
Algorithm 4 Entity Frequency Amplification (EFA): Algorithm for amplifying entity frequency via repeated insertion.
@@
Stopword Removal (SR): We remove stopwords from the context. Stopwords are commonly occurring words (viz. "the", "are", "to", etc.) that are assumed to have limited context specific information. Removal of such words should ideally not alter the response of an LLM if the LLM is certain of the answer. An example of this is seen in Table 1 fifth row, where the stopwords are striked out. We ensured that the negative words were not removed as they would change the meaning of the sentence.
$\begin{array}[]{rl}input&\;\leftarrow\;\texttt{remove stopwords from the inputsentence}\end{array}$
Algorithm 5 Stopword Removal (SR): Algorithm for removing stopwords from a document.
@@
Split Response Consistency (SRC): In this case like in the SD case the prompt is not perturbed. Rather the output is analyzed where it is randomly split such that each part is at least a single sentence. Semantic inconsistency between the two parts is measured using an NLI model’s (specifically, deberta-large-nli) contradiction probability, where one part is taken as the premise and the other the hypothesis. An example of this is seen in Table 1 last row, where the two sentences are clearly at odds with each other. This strategy though requires that the response is at least a couple of sentences long, thus responses with fewer than 3 sentences are removed.
$\begin{array}[]{rl}generated\_sentences&\;\leftarrow\;\texttt{Obtain one generation from the model using nucleus\_sampling}\\ sentences&\;\leftarrow\;\texttt{split the output into atomic sentences}\\ probability\_scores&\;\leftarrow\;[]\\ \texttt{Repeat 5 times}:\\ index&\;\leftarrow\;\texttt{Split output into two parts}\\ probability&\;\leftarrow\;\texttt{obtain the probability that chunk2 entails chunks1}\\ &\texttt{probability\_scores.append(probability)}\end{array}$
Algorithm 6 Split Response Consistency (SRC): Algorithm for checking consistency over random splits of a generated response.
@@
As seen in Table 9 in the appendix, the four perturbations above (PP, SP, EFA and SR) that alter the original prompt still maintain the semantics as intended in almost all cases.
$features\leftarrow[]$
$labels\leftarrow[]$
for d in datapoints do
       $feature\leftarrow[]$
       for p in perturbations do
             pert = perturb datapoint using perturbation p
             pert_generations = obtain model’s response for p
             semantic_sets = Set()
             setNum = 0
             Compute SRC Minimum Value
             if perturbation == SRC then
                   feature.append(max(probScores))
             end if
            for gen1, gen2 in pert_generations do
                   Compute number of semantic sets
                   if gen1 entails gens2 then
                         semanticSetNum = max(semantic set number of gen1, gen2 and 0)
                         if semanticSetNum == 0 then
                               semanticSetNum $\leftarrow$ setNum++
                               semantic_sets $\leftarrow$ assign semanticSetNum as the semantic number for gen1 and gen2
                         end if
                   end if
                  else
                         if gen1 not in semantic_sets then
                               semantic_sets $\leftarrow$ assign setNum++ as the semantic number of gen1
                         end if
                        if gen2 not in semantic_sets then
                               semantic_sets $\leftarrow$ assign setNum++ as the semantic number of gen2
                         end if
                   end if
             end for
            feature.append(number of semantic values)
             feature.append(rougeScore(pertGenerations))
       end for
      Compute label using lexical similarity
       label_generations $\leftarrow$ prompt the model to get 5 generations for d using nucleus sampling
       rs $\leftarrow$ compute rouge score of label_generations
       if rS > 0.3 then
             label = 1
       end if
      else
             label = 0
       end if
      labels.append(label)
end for
Algorithm 7 Pseudocode for generating features and labels
@@ 3.2 Featurization
SRC Minimum Value: As mentioned above, semantic inconsistency between the two parts is measured using an NLI models contradiction probability, where one part is taken as the premise and the other the hypothesis. The highest contradiction probability amongst multiple such partitions is the feature value for this strategy. In Table 1 last row, there are only two sentences so only one split would be done and since the sentences contradict each other the NLI contradiction probability would be high or consistency would be low. Note that optionally one can also pass the entire prompt as a feature in addition to the above. In the experiments, we saw minimal improvement with such an addition.
Semantic set and lexical similarity were first used by [18] where they applied it only for SD perturbation discussed previously. We describe the psuedocode of the entire end to end algorithm in 7
@@ 3.3 Label Creation and Confidence Estimation
4 Experiment Details
4.1 Models
We demonstrate the efficacy of our method on question answering and summarization tasks. For summarization, we used BART-large [22] and Pegasus-large [50] and for question answering, we used GPT-4 [30], Mistral-7B-Instruct-v0.2 [13], Llama-2-13b chat version [43], and Flan-ul2 models [40]. For detecting entailment, we use deberta-large-nli model which is specialized for NLI tasks [10].
4.2 Datasets
For question answering we elicited responses from these models on four datasets, namely, CoQA [36], SQuAD [34], TriviaQA [16] and Natural Questions (NQ) [20]. CoQA and SQuAD provide the context and expect the model to respond to the question based on the context, while TriviaQA and NQ do not have a context and require the model to tap into its learnt knowledge. For our experiments, we use the validation splits for all the datasets as done previously [24]. CoQA has 7983 datapoints, TriviaQA has 9960 datapoints, SQuAD has 10,600 datapoints and NQ has 7830 datapoints. For summarization, we used CNN  Daily Mail [37] and [11] and XSUM [29] datasets. We use a subset of the validation splits of both the datasets comprising of 4000 datapoints.
We follow previous works [24], which used $1000$ datapoints for hyperparameter tuning, to train our Logistic Regression Classifier and the rest of them were used for evaluation. For each of the prompt perturbations specified above, we use five generations for each perturbation for more robust evaluation. All results are averaged over five runs and we report standard deviations rounded to three decimal places for our method. We use zero-shot prompting for the datasets with context. For TriviaQA, Flan-ul2, Mistral-7B-Instruct-v0.2 and GPT-4 also worked well with zero shot prompting while Llama-2-13b chat was performant with a two-shot prompt. For NQ, we used a five shot prompt. The details about the prompts used are provided in the Appendix A.
@@ 4.2 Datasets
4.3 Compute
We used internally hosted models to generate the responses. Thus, we used V100s GPUs for the feature extraction step once the responses were generated. The logistic regression model was trained on an intel core CPU.
4.4 Baselines
5 Results
5.1 Confidence Estimation
Dataset Source LLM Self Target LLM 1 Target LLM 2 Target LLM 3
TriviaQA Llama 0.88 0.94 (Flan-ul2) 0.80 (Mistral) 0.94 (GPT-4)
Flan-ul2 0.94 0.87 (Llama) 0.80 (Mistral) 0.95 (GPT-4)
Mistral 0.81 0.84 (Llama) 0.91 (Flan-ul2) 0.95 (GPT-4)
GPT-4 0.96 0.85 (Llama) 0.92 (Flan-ul2) 0.80 (Mistral)
SQuAD Llama 0.83 0.81 (Flan-ul2) 0.80 (Mistral) 0.9 (GPT-4)
Flan-ul2 0.8 0.79 (Llama) 0.78 (Mistral) 0.89 (GPT-4)
Mistral 0.84 0.82 (Llama) 0.83 (Flan-ul2) 0.91 (GPT-4)
GPT-4 0.91 0.81 (Llama) 0.8 (Flan-ul2) 0.81 (Mistral)
CoQA Llama 0.92 0.79 (Flan-ul2) 0.78 (Mistral) 0.92 (GPT-4)
Flan-ul2 0.87 0.87 (Llama) 0.81 (Mistral) 0.92 (GPT-4)
Mistral 0.81 0.88 (Llama) 0.86 (Flan-ul2) 0.93 (GPT-4)
GPT-4 0.95 0.9 (Llama) 0.86 (Flan-ul2) 0.8 (Mistral)
NQ Llama 0.85 0.91 (Flan-ul2) 0.83 (Mistral) 0.91 (GPT-4)
Flan-ul2 0.93 0.83 (Llama) 0.82 (Mistral) 0.92 (GPT-4)
Mistral 0.83 0.85 (Llama) 0.90 (Flan-ul2) 0.91 (GPT-4)
GPT-4 0.93 0.84 (Llama) 0.91 (Flan-ul2) 0.81 (Mistral)
CNN Pegasus 0.77 0.57 (BART) - -
BART 0.57 0.77 (Pegasus) - -
XSUM Pegasus 0.73 0.58 (BART) - -
BART 0.57 0.71 (Pegasus) - -
Dataset Source LLM Self Target LLM 1 Target LLM 2 Target LLM 3
TriviaQA Llama 0.83 0.74 (Flan-ul2) 0.64 (Mistral) 0.85 (GPT-4)
Flan-ul2 0.74 0.83 (Llama) 0.64 (Mistral) 0.83 (GPT-4)
Mistral 0.64 0.83 (Llama) 0.73 (Flan-ul2) 0.86 (GPT-4)
GPT-4 0.89 0.81 (Llama) 0.71 (Flan-ul2) 0.63 (Mistral)
SQuAD Llama 0.68 0.62 (Flan-ul2) 0.63 (Mistral) 0.81 (GPT-4)
Flan-ul2 0.96 0.89 (Llama) 0.91 (Mistral) 0.82 (GPT-4)
Mistral 0.96 0.90 (Llama) 0.91 (Flan-ul2) 0.81 (GPT-4)
GPT-4 0.83 0.65 (Llama) 0.93 (Flan-ul2) 0.94 (Mistral)
CoQA Llama 0.71 0.79 (Flan-ul2) 0.61 (Mistral) 0.82 (GPT-4)
Flan-ul2 0.80 0.70 (Llama) 0.61 (Mistral) 0.81 (GPT-4)
Mistral 0.61 0.69 (Llama) 0.79 (Flan-ul2) 0.83 (GPT-4)
GPT-4 0.86 0.69 (Llama) 0.79 (Flan-ul2) 0.6 (Mistral)
NQ Llama 0.45 0.46 (Flan-ul2) 0.42 (Mistral) 0.77 (GPT-4)
Flan-ul2 0.47 0.45 (Llama) 0.42 (Mistral) 0.76 (GPT-4)
Mistral 0.42 0.45 (Llama) 0.46 (Flan-ul2) 0.77 (GPT-4)
GPT-4 0.79 0.43 (Llama) 0.46 (Flan-ul2) 0.41 (Mistral)
CNN Pegasus 0.74 0.34 (BART) - -
BART 0.34 0.74 (Pegasus) - -
XSUM Pegasus 0.27 0.34 (BART) - -
BART 0.35 0.25 (Pegasus) - -
Sample Efficiency Though we used 1000 datapoints to train the logistic regression classification, in Table 8, we show that our method is quite performant even with fewer training datapoints. This observation is also consistent when our method uses the same number of queries as our baselines as reported in 11, 12.
5.2 Confidence Model Interpretability and Transferability
Interpretability: We now study which features in our logistic model were instrumental for accurate confidence estimation. In Table 10 included in the appendix, we see the top four features for each dataset-LLM combination. Blanks indicate that there were no features at that rank or lower where their logistic coefficient was greater than $1e^{-4}$. As can be seen the simplest feature SD plays a role in many cases. This indicates that variability of output for the same input prompt is a strong indicator of response correctness. Moreover, other features such as SP and EFA are also crucial in ascertaining confidence as seen in particular for the SQuAD dataset as well as the summarization datasets. This points to order bias when looking at contexts and brittleness to redundant information being also strong indicators of response accuracy. PP and SR also play a role in some cases, where they are more crucial for datasets with no contexts such as TriviaQA and NQ. This makes sense as the specific question is more important here in the absence of context and hence the absence of also other features such as SP and EFA. Both the lexical similarity and semantic set featurizations seem to be important in estimating confidence.
Transferability: Given the commonality between the important features across LLMs for a dataset we now try to test how well does our logistic confidence model for one LLM perform in estimating confidences of another LLM. As seen in Tables 4 and 5 our confidence models are actually quite transferable as they perform comparably or even sometimes better on the other LLMs than the LLM they were built for. This is particularly true for Mistral where, its confidence model performs better for the other two LLMs than itself even coming close in performance to their own confidence models in many cases.
6 Discussion
In summary, we have provided an extensible framework for black-box confidence estimation of LLM responses by proposing novel features that are indicative of response correctness. By building an interpretable logistic regression model based on these features we were able to obtain state-of-the-art performance in estimating confidence on six benchmark datasets (CoQA, SQuAD, NQ, TriviaQA, CNN Daily and XSUM) and using six powerful LLMs (Llama-2-13b-chat, Mistral-7B-Instruct-v0.2, Flan-ul2, GPT-4, Pegasus-large and BART-large). The interpretability of our confidence model aided in identifying features (viz. SD, SP, EFA,PP) that were instrumental in driving its performance for different LLM-dataset combinations. This led to the interesting realization that many of the features crucial for performance were shared across the confidence models of different LLMs for a dataset. We thus tested if the confidence models generalized across LLMs for a dataset and found that it indeed was the case leading to the interesting possibility of having an universal confidence model trained on just a single LLMs responses, but applied across many others.
We used BART and Pegasus for summarization as they are specialized for this task. We tested on Flan-ul2 to validate that our method works on encoder-decoder architectures. Given that our approach worked for GPT-4 we believe it is generic enough to extend to other powerful LLMs. This includes reasoning models that generate solutions based of on chain-of-thought (COT) traces as these models although on the surface seem more robust suffer from creating incorrect COTs in many cases and sometimes “over think” making mistakes on simple problems [35, 2]. When and how much to “think” is still an active research topic. Our perturbations could alter the COTs affecting the outputs of such models especially when they are uncertain.
Owing to the supervised nature of training the confidence model, one limitation of our approach is that at least some of the model’s generations must be close to the ground truth for us to obtain a reasonable confidence estimator. We used rouge to test accuracy of generations consistent with previous works, however, rouge, like also other NLP metrics, can be error prone. Nonetheless, for summarization we also report results with GPT-4 as-a-judge in Tables 13 14 and 15 in the appendix, which are qualitatively similar to those reported in the main paper. In terms of broader impact, our approach can be widely applied as it is simple and works with just black-box access to the LLM. However, our estimates although accurate can be imperfect and this should be taken into account when using our approach in high stakes applications involving LLMs. One should also be cognizant of adversaries aware of our features trying to induce misplaced trust in LLMs they create or prefer.
Appendix A Prompt Design
Prompts for TriviaQA:
    Q: {question}
    A:
    Q: In Scotland a bothy/bothie is a?
    A: House
    Q: Who is Posh Spice in the spice girls pop band?
    A: Victoria Beckham
    Q: {question}
    A:
Prompts for CoQA
    context: {context}
    Question: {question}
    Answer: [/INST]
Prompts for SQuAD
    context: {context}
    Question: {question}
    Answer: [/INST]
Prompts for NQ: For all the models we used the following prompt:
Here are 5 Example Question Answer pairs:
Question: who makes up the state council in russia
Answer: governors and presidents
Question: when does real time with bill maher come back
Answer: November 9, 2018
Question: where did the phrase american dream come from
Answer: the mystique regarding frontier life
Question: what do you call a group of eels
Answer: bed
Question: who wrote the score for mission impossible fallout
Answer: Lorne Balfe
Question: {question}
Answer:
Appendix B Results